• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name aregjan

    Log entry time 08:04:22 on January24,2002

    Entry number 80550

    keyword=Shift Summary

    12:00-3:00 Doing beam position calibrations, and bulls-eye harp scans

    3:10am Some of the 1-16 wires on v2 wire chamber start ringing.
    Controled access is requested.

    3:24am, MCC called, sth. needs to be fixed, will be down for "a little while" (e.g., at least ~ 1 hour).

    3:30-5:10am Two accesses were needed to go and check the wire chambers. It
    was noticed that the problem goes and comes back. Hence
    it was decided that the problem must have come from a bad contact.
    All the preamp cards for V2 were pulled out and the contacts were rubbed
    with alcohol. The problem didn't seen to come back for awhile.

    5:15am Again v2 is recording huge rates, hence the above didn't help.
    Another access has been requested.

    5:45am No access was performed.
    It was noticed that in the case when the problem comes back, it is
    possible to eliminate it by raising the threshold on BTM VDC (look
    in the window "LEFT VDC - DISC LVL" in HAC) to 8.0 Volts.
    Hence the following is advised: in the case of a problem of ringing wire
    one should raise the threshold to 8.0 Volts for *1 minute*, and then
    lower to back to 4Volts.

    5:52 Raised BTM threshold to 8, and then back to 4.
    6:07 Wire ringing. Took less than 15 minutes.

    6:10am BTM threshold raised to 8V, then lowered to 4.5V. Waiting
    for the next ring.
    7:05am Nothing is still ringing, so far so good.

    7:05 am A problem in HAC - left colimator info. not available, exactly "infm". It's
    not clear how to move to sieve slit.

    7:45 am Ed Folts arrived to check the colimators.

    Runs 1592, 1593, 1594 were optics runs with open collimator.
    See the previous halog entry.